michelle has $6.82 in change. The number of pennies, dimes, nickels and quarters are, in that order, consecutive even integers. How many coins does she have?
Answer by richard1234(7193)

Let x, x+2, x+4, x+6 be the number of pennies, dimes, nickels, and quarters respectively. Then,

x + 10(x+2) + 5(x+4) + 25(x+6) = 682 (you see how the algebraic equation is put together?)

x + (10x + 20) + (5x + 20) + (25x + 150) = 682

41x + 190 = 682

41x = 492 --> x = 12, so Michelle has 12 pennies, 14 dimes, 16 nickels, and 18 quarters.